### v2

# Learn to Teach: Sample-Efficient Privileged Learning for Humanoid Locomotion over Diverse Terrains

Abstract
Humanoid robots promise transformative capabilities for industrial and service applications. While recent advances in Reinforcement Learning (RL) yield impressive results in locomotion, manipulation, and navigation, the proposed methods typically require enormous simulation samples to account for real-world variability. This work proposes a novel one-stage training framework—Learn to Teach (L2T)—which unifies teacher and student policy learning. Our approach recycles simulator samples and synchronizes the learning trajectories through shared dynamics, significantly reducing sample complexities and training time while achieving state-of-the-art performance. Furthermore, we validate the RL variant (L2T-RL) through extensive simulations and hardware tests on the Digit robot, demonstrating zero-shot sim-to-real transfer and robust performance over 12+ challenging terrains without depth estimation modules. Experimental videos are available online  [^1] https://lidar-learn-to-teach.github.io.
I Introduction
Reinforcement Learning (RL) has revolutionized robotic control by tackling complex tasks such as dynamic locomotion [1, 2, 3]. Despite these achievements, policies trained in simulators often falter when deployed into the real world due to the inevitable simulation-to-reality gap [4]. Although domain randomization [5] is widely used to mitigate these discrepancies, it incurs significantly higher sample complexity as agents must explore extensive environmental variations.
Recently, teacher-student learning methods have demonstrated promising results by leveraging an expert teacher to guide students with restricted observation spaces [6, 7, 8]. However, the conventional two-stage training discards valuable teacher interactions with the environment and often suffers from mismatches between independently trained teachers and students. To address these issues, we propose Learn-to-Teach (L2T): a unified training framework that co-trains teacher and student agents in a single, interactive stage, where the student fully utilizes the collected samples.
To quantify L2T’s advantages, we implement L2T-RL, an RL variant, and benchmark its performance on humanoid locomotion tasks using the Digit robot in Isaac Lab—a state-of-the-art GPU-accelerated simulator [9]. Our results show that L2T-RL can achieve stable and superior performance compared to the conventional teacher-student learning paradigm, requiring $50\%$ fewer samples. Consequently, we deploy our trained policy on the robot Digit and conduct extensive hardware experiments in indoor and outdoor environments. Strikingly, the resulting student agent, a lightweight LSTM-based policy, exhibits zero-shot sim2real transfer on the physical Digit robot across a wide range of terrains, including gravel, sand, grass, and slopes (Fig. 1). We also test our control policy on various perturbations such as push recovery, walking under payload, and walking on slippery or wet terrains or with the wind blowing (see Fig. 6 and the supplementary video). Our contributions are as follows:
Efficient training framework: We propose a joint teacher-student training paradigm that optimizes both policies simultaneously. Unlike prior decoupled approaches, our framework enables cross-agent knowledge transfer to the student policy by dynamically utilizing the teacher’s training samples directly within a single training stage, avoiding the need for training from scratch in a separate stage.
Mitigation of teacher-student policy divergence: We propose a sample mixing strategy to alleviate the imitation gap between the teacher and student, which traditional privileged learning is unable to address [10]. Both agents will contribute to the replay buffer following a predefined schedule when collecting samples. Mixing samples enables a joint optimization process that mitigates policy divergence while promoting sample efficiency by letting both agents explore Out-of-Distribution (OOD) data.
Humanoid RL agent deployment: We demonstrate real-world locomotion agility through hardware experiments. Our policy, trained entirely in simulation, enables a physical humanoid robot to reliably traverse $12+$ real-world terrains (concrete, gravel, slopes, stairs, etc.) and withstand dynamic perturbations (pushes and payloads) without offline fine-tuning. The policy achieves high success in unstructured environments, matching the teacher’s robustness despite using only proprioceptive inputs without depth estimation modules.
II Related work
Teacher-student learning: In the robotics learning community, teacher-student learning [6, 7, 8] has gained significant attention due to its applicability and effectiveness in addressing sim2real challenges. In this framework, the teacher agent is trained with complete knowledge of the state space. After obtaining an expert-level teacher, a student agent is trained in an observation space that follows the available sensor configurations on hardware, where the goal is to imitate the teacher’s action [8]. In this work, we extend this learning framework by training the teacher and the student simultaneously in a single stage. Prior work [11] proposed a method termed Concurrent Teacher Student (CTS) learning, which also explored the idea of co-training both agents. However, CTS trains a shared policy across the teacher and the student, only differentiating the observation encoder and the critic. This potentially disrupts the training process as the privileged critic naturally rewards actions that might not seem valuable to the student. In comparison, L2T trains two separate agents, with the option of sharing an encoder network or using an asymmetric learning style critic.
Figure 1: We implement our L2T-RL algorithm on our bipedal walking robot Digit and deploy it on diverse terrain with various environmental conditions such as wet grass, gravel, sandy terrain, and slippery surfaces.
Learning with partial observation: Recent advancements in RL under partial observability have significantly improved the ability of robotic systems to operate in complex, uncertain environments [8]. Contemporary approaches often leverage deep recurrent architectures, such as [12], to infer latent state representations from sequential data, effectively bridging traditional POMDP solvers with modern deep RL frameworks. In robotics control, practitioners construct history-dependent policies from a sliding-window style observation or rely on the recurrent architecture of the policy network. At the same time, asymmetric learning has emerged as another effective strategy to bridge the gap between training and execution [13]. In these approaches, the critic network is provided access to privileged, full-state information during training. Recent works have demonstrated that such asymmetric actor-critic frameworks improve sample efficiency and enhance policy robustness [14]. In this work, we combine these learning techniques, utilizing a recurrent network and an asymmetric critic, to solve the underlying POMDP problem efficiently.
Learning from demonstrations: Learning from demonstrations (LfD) has attracted significant interest in the robot learning field due to the growing abundance of robot data and the popularity of simple yet effective imitation learning (IL) frameworks [15]. LfD has demonstrated impressive results in controlling robot manipulators for tabletop tasks [16]. A recent surge of LfD studies in humanoids and bipeds have shown the promising potential of whole-body control and loco-manipulation [17, 18, 19, 16]. However, supervised learning demands high-quality behavior data, oftentimes through elaborate data collection pipelines [19, 16] and/or needs accurate re-targeting to robot states from datasets with different morphologies , e.g., internet video of human walking. The latter shares growing popularity due to its promising direction towards a generalist robotic control. However, supervised learning can fail to learn reliable whole-body control policies, so they depend on additional model-based balancing controllers [17]. On the other hand, the prevalent IL loss is known to be suboptimal from a pure learning perspective in previous studies [20]. Thus, in this work, we focus on a generic algorithm framework to address the sim2real gap alone, without the interference of possible issues brought up by LfD methods. Furthermore, our proposed framework can be easily extended to the LfD setting, which we leave as a future direction.
Figure 2: Learn to Teach (L2T) training pipeline. The teacher agent utilizes a neural network for the policy, which comprises three fully connected layers with sizes $[512,256,128]$. The student agent’s policy network deployed on the robot is an LSTM network with a hidden layer of $128$ units, followed by fully connected layers with shape $[512,256,128]$. The teacher learns via conventional RL methods, while the student updates its policy by imitating the teacher.
Bipedal locomotion over complex terrain: Humanoid robots recently have gained increasing interest due to their applicability and versatility [18, 21, 22, 17, 23, 24, 19, 25, 26, 27, 28], ranging from locomotion [29], to manipulation [30]. Prior bipedal locomotion works [29, 31] have explored the conventional teacher-student learning paradigm in locomotion tasks. However, the training process can take significant samples even with an elaborate training environment design. For instance, in [29], the teacher and the student agent require $1.7e9$ samples. In sharp contrast, we only require $3e8$ samples to train the same Digit robot to walk over challenging terrains for the same control task with even more terrain variations. Concurrent works also incorporate memory structure into the policy architecture [11, 29, 32], or learning from demonstrations collected from various data sources such as human motion data [33] or data generation using model-based methods [30]. In comparison, we design straightforward and intuitive reward functions for bipeds in general terrain settings, offering a simple yet effective solution.
III Methods
This section introduces our problem setup and notations and presents our learning framework. A Markov Decision Process (MDP), denoted as $\mathcal{M}$, is described by a tuple: $\mathcal{M}=\langle\mathcal{S},\mathcal{A},R,\mathcal{P},\Pi,\gamma\rangle$, where an agent starts with a given state $s_{0}$ following the initial state distribution $p(s_{0})$. At any time step $t$, the agent at the current state $s_{t}\in\mathcal{S}$ takes an action $a_{t}\in\mathcal{A}$ following the agent’s policy $\pi\in\Pi$, which defines a probability distribution over action space for each state. While receiving an instantaneous scalar reward $r(s_{t},a_{t})\in\mathbb{R}$, the state of the agent transitions to a new state $s_{t+1}\in\mathcal{S}$ following a transition model $\mathcal{P}(\cdot|s_{t},a_{t})$. $\gamma$ specifies the discount factor. The goal of the agent is to maximize the expected discounted sum of rewards the agent receives over time $\max_{\pi}\mathbb{E}\left[\sum^{\infty}_{t=0}\gamma^{t}r(s_{t},a_{t})\right]$, where the expectation is taken over actions $a_{t}\sim\pi(\cdot|s_{t})$, and transition probabilities $s_{t+1}\sim\mathcal{P}(\cdot|s_{t},a_{t})$ and initial distribution $s_{0}\sim p(s_{0})$. A Partially Observable Markov Decision Process (POMDP) is further coupled with an observation model $O(\cdot|s_{t})$, which is generally hidden from the agent. At each time step, the agent only observes $o_{t}\sim O(\cdot|s_{t})$ sampled from the observation model. Then, the agent takes an action based on $o_{t}$, following its policy $\pi(\cdot|o_{t})$, and subsequently receives a reward from the environment $r(s_{t},a_{t})$. To train a policy robust to various observation models, domain randomization is often applied. For example, by adding noise to the state $s_{t}$, the trained agent’s policy can handle observations $o_{t}=s_{t}+\epsilon$, where $\epsilon$ can be any noise distribution.
We train the teacher with a generic actor-critic method. During training, the teacher interacts with the environment, generates samples, and stores them in a replay buffer [34]. In standard teacher-student frameworks, the teacher’s collected samples are used solely for training the teacher policy $\pi_{t}$ and then discarded. In contrast, our L2T framework co-trains the student with the teacher, reusing the teacher’s samples across all iterations. Fig. 2 illustrates our learning framework, and Algorithm 1 presents the pseudo-code. Specifically, as the teacher interacts with the environment, we record samples $(s,a,r,s^{\prime})$, where $s^{\prime}$ denotes the next observation, the corresponding noisy observations $o$ generated by domain randomization, and $o^{\prime}$, the next observation. In other words, we store $(s,o,a,r,s^{\prime},o^{\prime})$ as training data in the replay buffer. The student updates its policy at each iteration by sampling mini-batches from the replay buffer, but its policy relies solely on the collected noisy data. This joint training procedure greatly improves sample efficiency as both agents learn together, without the need for a separate student training stage as that used in the traditional set-up.
Another key challenge in the teacher-student framework is the discrepancy between the teacher’s and the student’s observation spaces. Traditional teacher-student learning methods fall short because the teacher does not account for the limitations of the student’s observations, leading to suboptimal guidance [10]. To bridge this gap, we introduce a sample-mixing mechanism in which the student collects its own samples directly from the environment. These student-generated samples—including actions and the resulting observations—are incorporated into the replay buffer as if they were produced by the teacher. This injection of OOD data helps reduce the imitation gap between the two agents.
To systematically blend teacher and student experiences, we define a mixture coefficient $\alpha_{\text{mix}}$. At each time step, the action $a$ is determined by a probabilistic mixture of the teacher’s policy $\pi_{t}$ and the student’s policy $\pi_{s}$. Specifically, the action selection is defined as follows:
$a=\begin{cases}\text{sample }\pi_{s}(\cdot\mid o),&\text{with probability }\alpha_{\text{mix}},\\ \text{sample }\pi_{t}(\cdot\mid s),&\text{with probability }1-\alpha_{\text{mix}}.\end{cases}$ (1)
Additionally, $\alpha_{\text{mix}}$ is scheduled linearly from 0 to a predefined constant over the course of training, which is $0.2$ in our implementation. This formulation ensures that, initially, the teacher’s guidance dominates the action selection, but as training progresses, the student’s policy increasingly influences the learning. The scheduling of $\alpha_{\text{mix}}$ helps balance the contributions of both policies and guarantees the stability of the training process. By combining these strategies, our framework leverages both the teacher’s guidance and the student’s explorative capabilities to achieve more robust learning outcomes.
We implement a variant of our framework, L2T-RL. We apply policy gradient methods to update the teacher policy $\pi_{t}$. At each iteration $k$, the critic is updated by estimating the value functions: $V^{\pi}(s):=\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}\,r(s_{t},a_{t})\,|\,s_{0}=s\right],\\ Q^{\pi}(s,a):=\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}\,r(s_{t},a_{t})\,|\,s_{0}=s,a_{0}=a\right],$ where $V^{\pi}(s)$ is the value function, and $Q^{\pi}(s,a)$ is the discounted action-value function. For brevity, we will only use the subscript $t$ to denote the teacher from now on and use subscript $s$ to denote the student. Subsequently, the teacher’s policy is updated via a Policy Mirror Descent [35] step with a step size $\beta$:
$\min_{p_{t}}\;-\beta\langle Q^{\pi_{t}}(s,\cdot),p_{t}(\cdot|s)\rangle+\text{KL}(\pi_{t}\|p_{t})\quad\forall s\in\mathcal{S},$ (2)
where the optimal $p_{t}$ represents the teacher policy in the next iteration. Any policy improvement scheme can be fit into the L2T framework. This formulation encompasses a range of policy gradient methods [35], such as Proximal Policy Optimization (PPO) [36] and Soft Actor-Critic (SAC) [37]. In our experiments on the Digit robot, we employ a PPO-style update. However, our framework can be easily extended to various learning methods, including imitation learning methods, or Inverse Reinforcement Learning (IRL) methods such as in [20].
For the student policy, we consider two choices for loss functions. First, we can minimize an imitation loss between the teacher’s and the student’s policies:
$\min_{p_{s}}\;L_{IL}=\mathbb{E}_{s,\,o\sim D}\left\|p_{s}(\cdot|o)-\pi_{t}(\cdot|s)\right\|_{2},$ (3)
where $D$ denotes the replay buffer and the optimal $p_{s}$ represents the student policy in the next iteration. Alternatively, one may minimize the KL divergence between the two:
$\min_{p_{s}}\;L_{KL}=\mathbb{E}_{s,\,\sim D}\;\text{KL}\left(p_{s}(\cdot|o)\;\|\;\pi_{t}(\cdot|s)\right),$ (4)
or any statistical distance metric that fits the action space.
Besides the imitation loss, the student can be updated using an asymmetric learning approach [13] that leverages the teacher’s critic, i.e., the value functions:
$\min_{p_{s}}L_{\text{Asym}}=-\beta\langle Q^{\pi_{t}}(s,\cdot),p_{s}(\cdot|o)\rangle+\text{KL}(p_{s}\|\pi_{t}).$ (5)
We denote the general loss function for student agents as $L_{s}$. In our application on the Digit robot, we observed that using the $L_{2}$ imitation loss yields the best performance, while the addition of $L_{\text{Asym}}$ does not affect the overall performance by a large margin. We conjecture that the $L_{2}$ loss allows the student policy to have a slightly higher exploration capability as we observe that $L_{\text{Asym}}+L_{2}$ will reach a training plateau that is inferior in performance than using $L_{2}$ alone.
Algorithm 1 Learn to Teach - RL (L2T-RL)
1:Input: initial teacher policy $\pi_{t}^{0}$, student policy $\pi_{s}^{0}$, and step size sequences $\{\beta^{k}_{t}\}$ and $\{\beta^{k}_{s}\}$.
2:for $k=0$ to $K$ do
3:  Sample a mini-batch $D_{k}$ from replay buffer $D$.
4:  Update the teacher critic by estimating:
$Q^{\pi_{t}^{k+1}}(s,a),V^{\pi_{t}^{k+1}}(s).$
5:  Update the teacher policy:
$\pi^{k+1}_{t}=\arg\min_{p_{t}}-\beta^{k}_{t}\langle Q^{\pi_{t}^{k}}(s,\cdot),p_{t}(\cdot|s)\rangle+\text{KL}(\pi_{t}^{k}\|p_{t}).$
6:  Update the student policy:
$\pi^{k+1}_{s}=\arg\min_{\pi_{s}}L_{s}(\pi_{s}^{k}).$
7:  Roll out to collect new samples $D^{\prime}$ according to the scheduling in Eq. 1.
8:  Update the replay buffer: $D\leftarrow D\cup D^{\prime}$.
9:end for
IV Environment Design
The Digit robot is a bipedal walking robot with $30$ degrees of freedom, which includes $20$ actuated joints with $4$ per arm and $6$ per leg. All joints are revolute joints except for the shin and heel joints, which are spring-based. Notably, the Digit robot features three closed kinematic chains per leg. Two of these chains involve motors controlling the foot, assisted by additional rods, while the third chain is responsible for controlling the heel via a rod extending from the hip. This leg design makes it a significantly challenging task for RL algorithms due to the high-dimensional action space and the complex dynamics of the robot.
We highlight a significant portion of our work is to reconstruct a faithful Universal Scene Description (USD) model of the robot in IsaacLab [9], although this is not claimed as an algorithmic contribution. As a result, we build a velocity-tracking RL task with accurate dynamics w.r.t the robot hardware, facilitating the sim2real transfer.
IV-A Observation space
The observation space (see table I) is constructed using data provided by the robot’s sensors, including base linear velocity, base angular velocity, joint positions, and joint velocities. Additionally, we include the commanded velocity that the robot will receive from an external controller during execution and the computed projected gravity and gait phase based on the IMU data and the robot execution time. Finally, the actions in a previous time step are also incorporated into the observation space, which allows us to learn a history-dependent policy using recurrent neural nets. We model measurement noise as $o_{t}=s_{t}+\alpha\epsilon$ where $\alpha$ is the scale, and $\epsilon$ is either Gaussian or uniform noise. We add this noise to the student’s observation space to mimic the hardware sensors while keeping the teacher’s observation noise-free, except for the ones that incorporate the student’s observations in order to alleviate the imitation gap, which is considered a common practice. Additionally, privileged information (lower half of Table I) is provided for the teacher for easier training.
TABLE I: Observation Terms for Teacher and Student
Observation Terms Dim Noise Student $\pi_{s}$ Teacher $\pi_{t}$
Clock 2
Base lin. vel. 3
Base ang. vel. 3
Projected gravity 3
Velocity command 3
Joint pos. 30
Joint vel. 30
Last action 20
Base lin. vel. 3
Base ang. vel. 3
Joint position 48
Joint velocity 48
Root state (w) 7
Base lin. vel. (w) 3
Base ang. vel. (w) 3
Base pos. (w) 3
Base quant. (w) 4
Env params 316
Height scan 187
IV-B Action space
The action space is designed as the target full-body joint positions $q_{\rm target}$, which a Proportional Derivative (PD) controller will aim to track during execution. At a frequency of $50$ Hz, the policy predicts the current targeted joint based on the current observation, and then at a higher frequency ($1$ kHz), the PD controller computes the torque $\tau$ as inputs to the motors to control the robot’s joints. The target velocity is set to zero, which is commonly employed in legged robot research. The PD gains are determined through empirical tuning to ensure stable joint control. We use a standard PD control law for computing the torque, i.e, $\tau=\text{K}_{p}(q_{\text{target}}-q)+\text{K}_{d}(\dot{q}_{\text{target}}-\dot{q})$, where $q$ represents the measured joint positions and $\dot{q}$ represents the measured joint velocities.
IV-C Reward functions
Our reward function design is summarized in Table II. We adopt some of the existing reward functions in IsaacLab across other velocity command tasks for bipeds and quadrupeds, including termination penalty, action rate, joint deviation, etc. In addition, we design specific reward functions for training the Digit robot, for which we highlight two of them. For implementation details, please refer to our code.
TABLE II: Reward Functions and Their Weights
Reward Weight Reward Weight
Termination penalty -200.0 Foot contact 2.0
Being alive 0.01 Track foot height 0.5
Action rate -0.015 Foot clearance 0.5
DOF velocity -5e-4 Track lin vel XY 0.5
Undesired contacts -1.0 Track ang vel Z 1.0
Flat orientation -10.0 Lin vel XY -2.0
Feet air time 0.25 Ang vel Z -0.1
Feet sliding -1.0 DOF torques -1.0e-5
DOF pos limits -0.5 DOF acc -2.5e-7
Joint deviation hip -5.0 Feet air time 0.125
Joint deviation arms -0.3 Joint deviation toes -0.1
Track Foot Height: We reward the agent for following a desired foot height trajectory, which is precomputed as a quintic polynomial.
$r_{\mathrm{foot-track}}$ $=\exp\{-\|h_{\text{foot\_traj}}-h_{\text{foot\_traj\_target}}\|_{2}\},$ (6)
$h_{\text{foot\_traj}}$ is the actual foot height and $h_{\text{foot\_traj\_target}}$ is the target foot height. We adjust the desired foot height based on the current CoM position to adapt to uneven terrains. Specifically, instead of tracking the absolute height of the foot, we track the relative distance of the foot and the CoM position to compensate for the terrain height, which is hard to obtain at run time. This reward seems to be the most important reward term, without which we are unable to train an agent robust to challenging terrains.
Foot Contact Matching:
$r_{\mathrm{contact}}=\begin{cases}c_{1},&\text{sign}(\phi(t))=\text{sign}(F_{\text{GRF}}>0)\\ -c_{2},&\text{otherwise.}\end{cases}$ (7)
where $\phi(t)=\sin(2\pi t/h)$, $h=0.68$ is the gait cycle duration, $c_{1},c_{2}\in\mathbb{R}^{+}$ are constants. We reward the agent when the foot contact matches the desired gait cycle. For example, if $\phi(t)>0$ (indicating the foot should be in contact with the ground) and $F_{\text{GRF}}>0$ (indicating the foot is actually in contact), we assign a positive reward $c_{1}$. Otherwise, a penalty $c_{2}$ is applied.
TABLE III: Event Terms for Domain Randomization
Name Name
Rand. friction coeff Add base mass
Rand. gravity Add external force
Rand. base location Push robot
Rand. robot joints
IV-D Domain randomization and curriculum
We dynamically change the observation space and perturb the physical dynamics of the environment in the hope of capturing the randomness and variation of the real world. We list all the domain randomization schemes used in the training in Table. III. In practice, we find that randomizing friction coefficients and adding external pushes can greatly improve the robustness of the trained policy.
Additionally, we adopt the curriculum training setup implemented by IsaacLab. We utilize the existing terrain map, which includes seven different terrains: flat ground, slopes, stepping stones, and pyramid stairs up and down. Each environment, except for flat, has a difficulty parameter from 0 to $1$, where 0 stands for the easiest and $1$ the hardest. As an example, for the slope environment, the degree of the slope ranges from 0 to $40$. The environment is configured with a curriculum such that the agents’ performance will increase if they improve their tracking of the generated velocity command and decrease if they fail to do so.
TABLE IV: Neural Network Parameters
Parameter Value
Number of GPUs 1 RTX 4080
Isaac Lab Decimation 20
Simulation Time Step (dt) 0.001 s
Number of Environments 4096
Learning Epochs 5
Steps per Environment 12
Batch Size 256
Episode Length 1000
Discount Factor ($\gamma$) 0.99
Generalized Advantage Estimation ($\lambda$) 0.95
Entropy Coefficient 0.001
Clipping Parameter 0.2
Optimizer Adam
Learning Rate 5e-4
Learning Rate Schedule Linear Decay
Total Time Steps (Samples) 5e8
Number of Iterations (approx) 6100
Figure 3: Training curve against baselines. (a) We plot the difficulty of the terrain the agent can successfully traverse during training. (b) We train and evaluate each method on the flat ground and plot the velocity tracking of each agent.
V Computational Results
V-A Rough terrain difficulty progression
First, we show the progression of terrain difficulty (curriculum progression) throughout training, representing the overall task completion, i.e., the level of terrain difficulty the agent can walk over with the CoM velocity maintained within a specific range from the velocity command. This metric is calculated as the average difficulty level across the $6$ terrain setups mentioned earlier. In essence, faster learning corresponds to a steeper curve in terrain difficulty progression. Fig. 3(a) illustrates the terrain progression during training. The purple curve illustrates the training curve of L2T, reaching a maximum of $5.72$ level of difficulty. Compared to a two-stage traditional teacher-student learning (TSL) paradigm, which is represented by two curves: the teal curve represents the TSL teacher agent, and the blue curve represents the TSL student agent. Notice that in TSL, the student agent can only be trained after the teacher. The TSL teacher has an MLP-based policy, while the student is LSTM-based, sharing the same network architecture and hyperparameter as in L2T. It is noteworthy to mention that we can only obtain a reasonable training result by using data aggregation (DAGGER) [38], i.e., when training the student, periodically use the teacher to predict the action. Additionally, we include two RL methods for learning with student observation space using vanilla PPO, and recurrent PPO. It turns out that these two baselines perform significantly inferior. Since IsaacLab cannot be evaluated during training, the L2T curves represent the joint training progression between the L2T teacher and the student. One might think that the true performance of L2T students can be much worse than what the curve shows. However, as we observe that the L2T curve eventually goes higher than the TSL teacher, we argue that this curve is rather the performance lower bound for the L2T student.
V-B Flat terrain velocity tracking error
Next, we investigate the agent velocity tracking performance. Unfortunately velocity tracking is not a valid metric on general terrain due to the variability of the terrain difficulty. For example, the same agent might exhibit significantly different velocity tracking abilities either on slopes or stairs. To make a fair comparison, we re-train separate agents and collect the results on flat terrain, as it presents the most straightforward task for locomotion. We compare the velocity tracking between the CoM velocity and the commanded velocity on the $xy$ plane and the yaw angle. The results are plotted in Fig. 3(b). The L2T agent and the student from the traditional teacher-student learning (TSL) paradigm are comparable, with L2T better at yaw tracking and worse at tracking in the $xy$ plane. This shows that on the flat surface, L2T performs similarly to traditional teacher-student training paradigms but saves $50\%$ of the total samples. It is noteworthy to mention that we use the same network architecture and hyperparameters for both training frameworks. In Fig. 5(b)-(c), we also plot the recorded robot CoM sagittal velocity versus the commanded velocity and the robot’s CoM trajectory in the horizontal plane.
Figure 4: Left: in conventional TSL, the teacher supervises student. Right: L2T co-trains both agents together, where the student provides feedback to the teacher through mixed samples.
Figure 5: (a) Ablation study on the mixture coefficient $\alpha_{\rm mix}$. (b) CoM velocity versus commanded velocity. (c) xyxy trajectory with $x$ direction command.
Figure 6: (a) we perturb the robot with a harness. (b) we test the step-in-place motion on a crate of gravel. We let the robot walk over (c) a slippery terrain with (left) the Agility company controller and (right) our L2T controller. (d) rocky terrain with slopes. (e) elevated platforms. (f) hill with gravel.
V-C Mitigation of policy divergence
We observe that mixing samples can alleviate the imitation gap, which is caused by the teacher having access to privileged information that is unavailable to the student. This privileged information is marginalized during imitation learning for the student agent, resulting in the student agent requiring more exploration and acting more conservatively. For example, since the student does not know if it is at the edge of the stairs, they will act less confidently when walking downstairs. However, a trained expert teacher agent might act more confidently as it knows the structure of the stairs due to the privileged information of a local depth map. This is generally true due to the existence of observational noise. In the presence of the imitation gap, the teacher agent might generate desirable demonstrations only for the teacher itself, but not necessarily desirable ones for the student agent [10].
In traditional TSL paradigms, this problem can be partially addressed by DAGGER or asymmetric learning. In L2T, all these techniques are naturally blended together as a single-stage co-training process, providing an efficient and elegant solution to address the imitation gap and policy divergence issue. We show that in Fig. 4, the walking gait of the student trained by DAGGER is significantly different from the teacher policy, while L2T can faithfully imitate the teacher. Notice the shape of the toe pad from the L2T student, which slightly tilts up from the horizontal plane, accurately mimicking the teacher’s toe. In contrast, the DAGGER-trained student agent has a flat-ground walking gait, with the toe pad parallel to the horizontal plane.
V-D Ablation study on the mixture coefficient
We conduct an ablation study on the critical components to further explore key design choices within our algorithm framework. Fig. 5(a) demonstrates the importance of the mixture coefficient $\alpha_{\text{mix}}$ in the training process. We observe that while the algorithm becomes unstable with a large $\alpha_{\text{mix}}$, which is caused by letting a highly suboptimal student agent inject too many samples, an appropriate chosen $\alpha_{\text{mix}}$ can benefit the overall training process. We hypothesize that the discrepancies between the student and teacher promote exploration within the action space, enabling both agents to learn from a broader region around the teacher’s actions.
VI Hardware Experiments
VI-A Locomotion over challenging terrains
We report the results of deploying our policy in various outdoor environments, as illustrated in Fig. 1 and the supplementary video. These experiments were conducted around a university campus, including pedestrian walkways, wooden bridges, grass-covered hills, beach volleyball courts, and gravel paths. Extensive outdoor testing highlights the policy’s capability to maintain a reliable and natural walking gait and withstand diverse environmental conditions, such as walking on wet surfaces and operating with strong wind.
Surprisingly, the policy shows generalization ability to scenarios not included in the training. For example, on grass hills, due to rain, the grass and the soil underneath exhibit a certain level of deformation upon impact, which increases the difficulty of state estimation and thus further increases the observation noise. Moreover, the policy can walk on the beach volleyball court with sandy terrain, as shown in Fig. 1. Despite the robot not being calibrated or trained for such conditions, the policy adapts to the environmental changes without additional training. This demonstrates the robustness and adaptability of our approach. Next, we examine the policy’s performance on terrains with obstacles. The first test involves a crate of gravel shown in Fig. 6(b). The robot consistently performs stepping-in-place actions with a stable motion. Even when the robot occasionally strikes the crate’s edge, it recovers and resumes normal a stepping gait. In the second test, we deploy the robot on a slippery terrain, where we distribute poppy seeds on a whiteboard shown in Fig. 6(c). Surprisingly, our robot does not exhibit any perceptible shaking motions. In contrast, the company controller provided by Agility Robotics fails to walk over. Furthermore, the policy enables omnidirectional control of the robot. Experiments with velocity commands demonstrate the robot’s directional control, which is crucial for practical deployment. Our policy can sustain some deformation and irregularity but may drift and lose balance on such occasions. Readers can find these experiments in the video.
VI-B Perturbation experiments
We evaluate the policy’s response to external perturbations applied during locomotion. Two scenarios are considered: pushing the robot’s center of mass (CoM) from the front, and back using a stick. The policy demonstrates robustness to withstand frontal and rearward pushes while maintaining a stable walking gait. We could not push the robot to failure even when applying a large force. Additionally, we apply a more substantial perturbation by using a harness to pull the robot with an impulsive force (see Fig. 6(a)). The robot adapts dynamically, exhibiting agile adjustments to its gait to compensate for the external perturbations.
VII Conclusion
We introduced L2T-RL, a novel single-stage learning framework that unifies teacher and student training to address sample inefficiency and enhance real-world performance. Our extensive simulation and hardware experiments demonstrate that L2T-RL achieves robust, agile, and precise locomotion while reducing sample complexity by $50\%$ and dramatically saving training time. These results highlight our contributions to redefining teacher-student learning paradigms and paving the way for practical RL-based robotic systems.

### v3

@@
-Learn to Teach: Sample-Efficient Privileged Learning for Humanoid Locomotion over Diverse Terrains
+Learn to Teach: Sample-Efficient Privileged Learning for Humanoid Locomotion over Real-World Uneven Terrain
 Abstract
@@
 Our approach recycles simulator samples and synchronizes the learning trajectories through shared dynamics, significantly reducing sample complexities and training time while achieving state-of-the-art performance.
-Furthermore, we validate the RL variant (L2T-RL) through extensive simulations and hardware tests on the Digit robot, demonstrating zero-shot sim-to-real transfer and robust performance over 12+ challenging terrains without depth estimation modules.
-Experimental videos are available online  [^1] https://lidar-learn-to-teach.github.io.
+Furthermore, we validate the RL variant (L2T-RL) through extensive simulations and hardware tests on the Digit robot, demonstrating zero-shot sim-to-real transfer and robust performance over 12+ diverse terrains without depth estimation modules.
+Experimental videos are available at https://lidar-learn-to-teach.github.io.
+Index Terms:
+Reinforcement Learning, Humanoid and Bipedal Locomotion, Sim2Real, Sample Efficiency
 I Introduction
@@
 Unlike prior decoupled approaches, our framework enables cross-agent knowledge transfer to the student policy by dynamically utilizing the teacher’s training samples directly within a single training stage, avoiding the need for training from scratch in a separate stage.
-Mitigation of teacher-student policy divergence: We propose a sample mixing strategy to alleviate the imitation gap between the teacher and student, which traditional privileged learning is unable to address [10].
+Mitigation of teacher-student imitation gap: We propose a sample mixing strategy to alleviate the imitation gap between the teacher and student, which traditional privileged learning is unable to address [10].
 Both agents will contribute to the replay buffer following a predefined schedule when collecting samples.
-Mixing samples enables a joint optimization process that mitigates policy divergence while promoting sample efficiency by letting both agents explore Out-of-Distribution (OOD) data.
+Mixing samples enables a joint optimization process that mitigates the imitation gap while promoting sample efficiency by letting both agents explore Out-of-Distribution (OOD) data.
 Humanoid RL agent deployment: We demonstrate real-world locomotion agility through hardware experiments.
@@
 A recent surge of LfD studies in humanoids and bipeds have shown the promising potential of whole-body control and loco-manipulation [17, 18, 19, 16].
-However, supervised learning demands high-quality behavior data, oftentimes through elaborate data collection pipelines [19, 16] and/or needs accurate re-targeting to robot states from datasets with different morphologies , e.g., internet video of human walking.
-The latter shares growing popularity due to its promising direction towards a generalist robotic control.
-However, supervised learning can fail to learn reliable whole-body control policies, so they depend on additional model-based balancing controllers [17].
-On the other hand, the prevalent IL loss is known to be suboptimal from a pure learning perspective in previous studies [20].
+However, supervised learning demands high-quality behavior data, oftentimes through elaborate data collection pipelines [19, 16] and/or needs accurate re-targeting to robot states from datasets with different morphologies On the other hand, the prevalent IL loss is known to be suboptimal from a learning perspective [20].
 Thus, in this work, we focus on a generic algorithm framework to address the sim2real gap alone, without the interference of possible issues brought up by LfD methods.
@@
 The teacher learns via conventional RL methods, while the student updates its policy by imitating the teacher.
-Bipedal locomotion over complex terrain: Humanoid robots recently have gained increasing interest due to their applicability and versatility [18, 21, 22, 17, 23, 24, 19, 25, 26, 27, 28], ranging from locomotion [29], to manipulation [30].
+Bipedal locomotion over complex terrain: Humanoid robots recently have gained increasing interest due to their applicability and versatility [18, 21, 22, 17, 23, 24, 19, 25], ranging from locomotion [26], to manipulation [27].
 Prior bipedal locomotion works [26, 28] have explored the conventional teacher-student learning paradigm in locomotion tasks.
 However, the training process can take significant samples even with an elaborate training environment design.
-For instance, in [29], the teacher and the student agent require $1.7e9$ samples.
-In sharp contrast, we only require $3e8$ samples to train the same Digit robot to walk over challenging terrains for the same control task with even more terrain variations.
-Concurrent works also incorporate memory structure into the policy architecture [11, 29, 32], or learning from demonstrations collected from various data sources such as human motion data [33] or data generation using model-based methods [30].
+Concurrent works also incorporate memory structure into the policy architecture [11, 26, 29], or learning from demonstrations collected from various data sources such as human motion [30] or generation using model-based methods [27].
 In comparison, we design straightforward and intuitive reward functions for bipeds in general terrain settings, offering a simple yet effective solution.
@@
 2 illustrates our learning framework, and Algorithm 1 presents the pseudo-code.
+We employ an MLP for the teacher because, given privileged access to the full state (depth scans, root pose, terrain profile, etc.), each observation is already fully descriptive and Markovian so that there is no need to model temporal dependencies via recurrence.
+We also notice that this architecture has been widely adopted by concurrent teacher-student learning paradigms [26].
 Specifically, as the teacher interacts with the environment, we record samples $(s,a,r,s^{\prime})$, where $s^{\prime}$ denotes the next observation, the corresponding noisy observations $o$ generated by domain randomization, and $o^{\prime}$, the next observation.
@@
 The student updates its policy at each iteration by sampling mini-batches from the replay buffer, but its policy relies solely on the collected noisy data.
-This joint training procedure greatly improves sample efficiency as both agents learn together, without the need for a separate student training stage as that used in the traditional set-up.
+This joint training procedure greatly improves sample efficiency as both agents learn together, without the need for a separate stage as used in the traditional set-up.
 Another key challenge in the teacher-student framework is the discrepancy between the teacher’s and the student’s observation spaces.
@@
 At each time step, the action $a$ is determined by a probabilistic mixture of the teacher’s policy $\pi_{t}$ and the student’s policy $\pi_{s}$.
+As such, we collect trajectories $s_{1},a_{1},s_{2},a_{2},s_{3},a_{3}$ where $a_{i}$ could either come from the teacher or the student.
+This sample mixing mechanism ensures that information from the student agent back propagates to the teacher.
 Specifically, the action selection is defined as follows:
@@
 This formulation ensures that, initially, the teacher’s guidance dominates the action selection, but as training progresses, the student’s policy increasingly influences the learning.
-The scheduling of $\alpha_{\text{mix}}$ helps balance the contributions of both policies and guarantees the stability of the training process.
+The scheduling of $\alpha_{\text{mix}}$ helps balance the contributions of both policies and guarantees stable training.
 By combining these strategies, our framework leverages both the teacher’s guidance and the student’s explorative capabilities to achieve more robust learning outcomes.
+This is in steep contrast to CTS, where the teacher encoder and the student encoder trains on independent datasets.
 We implement a variant of our framework, L2T-RL.
@@
 However, our framework can be easily extended to various learning methods, including imitation learning methods, or Inverse Reinforcement Learning (IRL) methods such as in [20].
+In practice, we pool samples from both agents to update the teacher policy using PPO.
+If we can maintain $\pi_{s}\approx\pi_{t}$, any student action $a_{s}\sim\pi_{s}(\cdot|s)$ satisfies $\pi_{t}(a_{s}|s)>0$, allowing us to view these student-generated transitions as valid (albeit lower-probability) samples from the teacher’s distribution.
 For the student policy, we consider two choices for loss functions. First, we can minimize an imitation loss between the teacher’s and the student’s policies:
-$\min_{p_{s}}\;L_{IL}=\mathbb{E}_{s,\,o\sim D}\left\|p_{s}(\cdot|o)-\pi_{t}(\cdot|s)\right\|_{2},$ (3)
+$\min_{p_{s}}\;L_{\text{IL}}=\mathbb{E}_{s,\,o\sim D}\left\|p_{s}(\cdot|o)-\pi_{t}(\cdot|s)\right\|_{2},$ (3)
 where $D$ denotes the replay buffer and the optimal $p_{s}$ represents the student policy in the next iteration. Alternatively, one may minimize the KL divergence between the two:
-$\min_{p_{s}}\;L_{KL}=\mathbb{E}_{s,\,\sim D}\;\text{KL}\left(p_{s}(\cdot|o)\;\|\;\pi_{t}(\cdot|s)\right),$ (4)
+$\min_{p_{s}}\;L_{KL}=\mathbb{E}_{s,\,o\sim D}\;\text{KL}\left(p_{s}(\cdot|o)\;\|\;\pi_{t}(\cdot|s)\right),$ (4)
 or any statistical distance metric that fits the action space.
 Besides the imitation loss, the student can be updated using an asymmetric learning approach [13] that leverages the teacher’s critic, i.e., the value functions:
-$\min_{p_{s}}L_{\text{Asym}}=-\beta\langle Q^{\pi_{t}}(s,\cdot),p_{s}(\cdot|o)\rangle+\text{KL}(p_{s}\|\pi_{t}).$ (5)
+$\min_{p_{s}}L_{\text{Asym}}=-\beta\langle Q^{\pi_{t}}(s,\cdot),p_{s}(\cdot|o)\rangle+\text{KL}(p_{s}\|\pi_{s}).$ (5)
 We denote the general loss function for student agents as $L_{s}$.
-In our application on the Digit robot, we observed that using the $L_{2}$ imitation loss yields the best performance, while the addition of $L_{\text{Asym}}$ does not affect the overall performance by a large margin.
-We conjecture that the $L_{2}$ loss allows the student policy to have a slightly higher exploration capability as we observe that $L_{\text{Asym}}+L_{2}$ will reach a training plateau that is inferior in performance than using $L_{2}$ alone.
+In our application on the Digit robot, we observed that using the $L_{\text{IL}}$ imitation loss yields the best performance, while the addition of $L_{\text{Asym}}$ does not affect the overall performance by a large margin.
+We conjecture that the $L_{\text{IL}}$ loss allows the student policy to have a slightly higher exploration capability as we observe that $L_{\text{Asym}}+L_{\text{IL}}$ will reach a training plateau that is inferior in performance than using $L_{\text{IL}}$ alone.
 Algorithm 1 Learn to Teach - RL (L2T-RL)
-1:Input: initial teacher policy $\pi_{t}^{0}$, student policy $\pi_{s}^{0}$, and step size sequences $\{\beta^{k}_{t}\}$ and $\{\beta^{k}_{s}\}$.
-2:for $k=0$ to $K$ do
-3:  Sample a mini-batch $D_{k}$ from replay buffer $D$.
-4:  Update the teacher critic by estimating:
-$Q^{\pi_{t}^{k+1}}(s,a),V^{\pi_{t}^{k+1}}(s).$
-5:  Update the teacher policy:
-$\pi^{k+1}_{t}=\arg\min_{p_{t}}-\beta^{k}_{t}\langle Q^{\pi_{t}^{k}}(s,\cdot),p_{t}(\cdot|s)\rangle+\text{KL}(\pi_{t}^{k}\|p_{t}).$
-6:  Update the student policy:
-$\pi^{k+1}_{s}=\arg\min_{\pi_{s}}L_{s}(\pi_{s}^{k}).$
-7:  Roll out to collect new samples $D^{\prime}$ according to the scheduling in Eq. 1.
-8:  Update the replay buffer: $D\leftarrow D\cup D^{\prime}$.
-9:end for
+0: initial teacher policy $\pi_{t}^{0}$, student policy $\pi_{s}^{0}$, and step size sequences $\{\beta^{k}_{t}\}$ and $\{\beta^{k}_{s}\}$
+1:for $k=0$ to $K$ do
+2:  Sample a mini-batch $D_{k}$ from the replay buffer $D$
+3:  Update the teacher critic $Q^{\pi_{t}^{k+1}}(s,a),V^{\pi_{t}^{k+1}}(s)$
+4:  Update the teacher policy:
+$\pi^{k+1}_{t}=\arg\min_{p_{t}}\left[-\beta^{k}_{t}\langle Q^{\pi_{t}^{k}}(s,\cdot),p_{t}(\cdot|s)\rangle+\mathrm{KL}(\pi_{t}^{k}\|p_{t})\right]$
+5:  Update the student policy:
+$\pi^{k+1}_{s}=\arg\min_{\pi_{s}}L_{s}(\pi_{s}^{k})$
+6:  Roll out to collect new samples $D^{\prime}$ according to the scheduling in Eq. 1
+7:  Update the replay buffer: $D\leftarrow D\cup D^{\prime}$
+8:end for
 IV Environment Design
@@
 We highlight a significant portion of our work is to reconstruct a faithful Universal Scene Description (USD) model of the robot in IsaacLab [9], although this is not claimed as an algorithmic contribution.
-As a result, we build a velocity-tracking RL task with accurate dynamics w.r.t the robot hardware, facilitating the sim2real transfer.
+As a result, we build a velocity-tracking RL task with accurate dynamics w.r.t the robot hardware.
 IV-A Observation space
 The observation space (see table I) is constructed using data provided by the robot’s sensors, including base linear velocity, base angular velocity, joint positions, and joint velocities.
-Additionally, we include the commanded velocity that the robot will receive from an external controller during execution and the computed projected gravity and gait phase based on the IMU data and the robot execution time.
+Additionally, we include the commanded velocity that the robot will receive from an external controller during execution, the computed projected gravity based on the IMU data, and desired gait phase based on the robot execution time.
 Finally, the actions in a previous time step are also incorporated into the observation space, which allows us to learn a history-dependent policy using recurrent neural nets.
@@
 Additionally, privileged information (lower half of Table I) is provided for the teacher for easier training.
+While our locomotion experiments randomize only the base mass, we include all environment parameters as privileged information, bundling them simplifies our environment API for future extensions.
+Note that the base linear velocity is given by the low-level software APIs provided by Agility Robotics.
 TABLE I: Observation Terms for Teacher and Student
 Observation Terms Dim Noise Student $\pi_{s}$ Teacher $\pi_{t}$
-Clock 2
+Clock input 2
 Base lin. vel. 3
@@
 Last action 20
-Base lin. vel. 3
-Base ang. vel. 3
-Joint position 48
-Joint velocity 48
 Root state (w) 7
@@
 DOF pos limits -0.5 DOF acc -2.5e-7
-Joint deviation hip -5.0 Feet air time 0.125
-Joint deviation arms -0.3 Joint deviation toes -0.1
+Joint deviation hip -5.0 Joint deviation toes -0.1
+Joint deviation arms -0.3
 Track Foot Height: We reward the agent for following a desired foot height trajectory, which is precomputed as a quintic polynomial.
@@
 Specifically, instead of tracking the absolute height of the foot, we track the relative distance of the foot and the CoM position to compensate for the terrain height, which is hard to obtain at run time.
-This reward seems to be the most important reward term, without which we are unable to train an agent robust to challenging terrains.
+This reward produces significantly different motions than rewarding foot clearance alone and it is one of the most important reward term, without which we are unable to train an agent robust to uneven terrains.
+Note that, this reward is different from the foot clearance reward as the latter rewards the foot to reach a desired height relative to floor.
+On flat ground, the desired height is consistent with the peak of the reference foot trajectory.
 Foot Contact Matching:
@@
 We utilize the existing terrain map, which includes seven different terrains: flat ground, slopes, stepping stones, and pyramid stairs up and down.
-Each environment, except for flat, has a difficulty parameter from 0 to $1$, where 0 stands for the easiest and $1$ the hardest.
-As an example, for the slope environment, the degree of the slope ranges from 0 to $40$.
-The environment is configured with a curriculum such that the agents’ performance will increase if they improve their tracking of the generated velocity command and decrease if they fail to do so.
-TABLE IV: Neural Network Parameters
-Parameter Value
-Number of GPUs 1 RTX 4080
-Isaac Lab Decimation 20
-Simulation Time Step (dt) 0.001 s
-Number of Environments 4096
-Learning Epochs 5
-Steps per Environment 12
-Batch Size 256
-Episode Length 1000
-Discount Factor ($\gamma$) 0.99
-Generalized Advantage Estimation ($\lambda$) 0.95
-Entropy Coefficient 0.001
-Clipping Parameter 0.2
-Optimizer Adam
-Learning Rate 5e-4
-Learning Rate Schedule Linear Decay
-Total Time Steps (Samples) 5e8
-Number of Iterations (approx) 6100
 Figure 3: Training curve against baselines.
-(a) We plot the difficulty of the terrain the agent can successfully traverse during training.
-(b) We train and evaluate each method on the flat ground and plot the velocity tracking of each agent.
+Left: We plot the difficulty of the terrain the agent can successfully traverse during training.
+Middle: we show the student training curve in TSL after obtaining a teacher.
+Right: average success rate of the student agents across $4096$ training environments with randomly sampled velocity commands.
 V Computational Results
@@
 Fig.
-3(a) illustrates the terrain progression during training.
-The purple curve illustrates the training curve of L2T, reaching a maximum of $5.72$ level of difficulty.
-Compared to a two-stage traditional teacher-student learning (TSL) paradigm, which is represented by two curves: the teal curve represents the TSL teacher agent, and the blue curve represents the TSL student agent.
+3 illustrates the terrain progression during training.
+We represent the teacher policy with a solid blue line and the student policy with a dashed blue line.
+We compare our approach against three baseline methods based on the asymmetric learning framework: CTS [11], ROA [35], and DWL [36].
+In addition, we include comparisons with the conventional teacher-student learning (TSL) paradigm and a vanilla recurrent PPO trained directly in the student’s environment.
 Notice that in TSL, the student agent can only be trained after the teacher.
@@
 It is noteworthy to mention that we can only obtain a reasonable training result by using data aggregation (DAgger) [37], i.e., when training the student, periodically use the teacher to predict the action.
-Additionally, we include two RL methods for learning with student observation space using vanilla PPO, and recurrent PPO.
-It turns out that these two baselines perform significantly inferior.
-Since IsaacLab cannot be evaluated during training, the L2T curves represent the joint training progression between the L2T teacher and the student.
-One might think that the true performance of L2T students can be much worse than what the curve shows.
-However, as we observe that the L2T curve eventually goes higher than the TSL teacher, we argue that this curve is rather the performance lower bound for the L2T student.
-V-B Flat terrain velocity tracking error
-Next, we investigate the agent velocity tracking performance.
-Unfortunately velocity tracking is not a valid metric on general terrain due to the variability of the terrain difficulty.
-For example, the same agent might exhibit significantly different velocity tracking abilities either on slopes or stairs.
-To make a fair comparison, we re-train separate agents and collect the results on flat terrain, as it presents the most straightforward task for locomotion.
-We compare the velocity tracking between the CoM velocity and the commanded velocity on the $xy$ plane and the yaw angle.
-The results are plotted in Fig.
-3(b).
-The L2T agent and the student from the traditional teacher-student learning (TSL) paradigm are comparable, with L2T better at yaw tracking and worse at tracking in the $xy$ plane.
-This shows that on the flat surface, L2T performs similarly to traditional teacher-student training paradigms but saves $50\%$ of the total samples.
-It is noteworthy to mention that we use the same network architecture and hyperparameters for both training frameworks.
-In Fig.
-5(b)-(c), we also plot the recorded robot CoM sagittal velocity versus the commanded velocity and the robot’s CoM trajectory in the horizontal plane.
-Figure 4: Left: in conventional TSL, the teacher supervises student. Right: L2T co-trains both agents together, where the student provides feedback to the teacher through mixed samples.
-Figure 5: (a) Ablation study on the mixture coefficient $\alpha_{\rm mix}$. (b) CoM velocity versus commanded velocity. (c) xyxy trajectory with $x$ direction command.
-Figure 6: (a) we perturb the robot with a harness.
-(b) we test the step-in-place motion on a crate of gravel.
-We let the robot walk over (c) a slippery terrain with (left) the Agility company controller and (right) our L2T controller.
-(d) rocky terrain with slopes.
-(e) elevated platforms.
-(f) hill with gravel.
-V-C Mitigation of policy divergence
+Discussion: First, Fig.
+3 shows that L2T significantly outperforms asymmetric learning baselines, including CTS, ROA, and DWL.
+We are unable to reproduce competitive performance with DWL due to missing implementation details.
+However, as DWL follows a similar framework to ROA, learning an encoder to reconstruct privileged states, we believe its potential for improvement over ROA is limited.
+Since these (single actor) methods lack a separate teacher policy, the student must discover high-reward states with a suboptimal student encoder in order to learn from them, which leads to poor sample efficiency.
+Furthermore, we observe a substantial imitation gap between the teacher and student in CTS.
+This arises because CTS processes teacher and student samples independently, relying solely on a reconstruction loss to align the two encoders.
+In contrast, L2T mixes samples from both agents, allowing the teacher to observe and evaluate the student’s behavior.
+From a data-driven viewpoint, CTS distills knowledge using two agents trained on disjoint datasets, while L2T actively injects out-of-distribution student trajectories into the teacher’s training, enabling more effective guidance of the student policy.
+Second, compared to the conventional teacher-student learning (TSL) paradigm, we find that L2T enables the student to achieve comparable performance without requiring a separate teacher pretraining stage.
+That is, the time spent training the teacher in TSL is sufficient for jointly training both teacher and student in L2T, making it a drop-in replacement.
+While we do not claim that L2T outperforms TSL in final policy quality, we save equivalently $50\%$ training time, or roughly $12$ hours of GPU time, a significant improvement in reducing development cycles.
+Ultimately, both approaches are upper-bounded by the performance of the teacher policy.
+(a) Imitation gap ablation.
+(b) Loss function ablation.
+Figure 4: Ablation studies on sample mixing and loss functions.
+V-B Mitigation of imitation gap
 We observe that mixing samples can alleviate the imitation gap, which is caused by the teacher having access to privileged information that is unavailable to the student.
@@
 In the presence of the imitation gap, the teacher agent might generate desirable demonstrations only for the teacher itself, but not necessarily desirable ones for the student agent [10].
-In traditional TSL paradigms, this problem can be partially addressed by DAGGER or asymmetric learning.
-In L2T, all these techniques are naturally blended together as a single-stage co-training process, providing an efficient and elegant solution to address the imitation gap and policy divergence issue.
+In L2T, imitation learning and sample mixing create a bidirectional feedback loop.
+The teacher policy benefits from student-generated samples, while the student continuously tracks the evolving teacher.
+This approach resembles techniques that augment training data [38] in supervised learning works such as computer vision, where randomization injects out-of-distribution data into the training set.
+In the context of our approach, the out-of-distribution data, with respect to the teacher policy, are from the student.
+Figure 5: Our method can mitigate the imitation gap, enabling accurately track teacher policy behavior.
 We show that in Fig.
-4, the walking gait of the student trained by DAGGER is significantly different from the teacher policy, while L2T can faithfully imitate the teacher.
+5, the walking gait of the student trained by DAgger is significantly different from the teacher policy, while L2T can faithfully imitate the teacher.
 Notice the shape of the toe pad from the L2T student, which slightly tilts up from the horizontal plane, accurately mimicking the teacher’s toe.
 In contrast, the DAgger-trained student agent has a flat-ground walking gait, with the toe pad parallel to the horizontal plane.
-V-D Ablation study on the mixture coefficient
+Details of the walking motions can be found in our video.
+Additionally, we compute the $L_{\text{IL}}$ between the teacher and student policies under three settings: L2T, TSL, and L2T without sample mixing ($\alpha_{\text{mix}}=0$) denoted as L2T w/ No Mix as shown in Fig.
+4a.
+This highlights the discrepancy between the two agents across methods and demonstrates that L2T effectively reduces the imitation gap.
+Figure 6: (a) Ablation study on the mixture coefficient. (b) CoM velocity versus command using student policy. (c) xyxy trajectory with $x$ direction command using student policy.
+V-C Ablation study
 We conduct an ablation study on the critical components to further explore key design choices within our algorithm framework.
 Fig.
-5(a) demonstrates the importance of the mixture coefficient $\alpha_{\text{mix}}$ in the training process.
+6(a) demonstrates the importance of the mixture coefficient $\alpha_{\text{mix}}$ in the training process.
 We observe that while the algorithm becomes unstable with a large $\alpha_{\text{mix}}$, which is caused by letting a highly suboptimal student agent inject too many samples, an appropriate chosen $\alpha_{\text{mix}}$ can benefit the overall training process.
 We hypothesize that the discrepancies between the student and teacher promote exploration within the action space, enabling both agents to learn from a broader region around the teacher’s actions.
+Additionally, we show in Fig.
+4b that choosing $L_{\text{IL}}$ slightly outperforms $L_{\text{IL}}+L_{\text{Asym}}$ in our training setup.
 VI Hardware Experiments
-VI-A Locomotion over challenging terrains
+VI-A Locomotion over real-world uneven terrain
 We report the results of deploying our policy in various outdoor environments, as illustrated in Fig.
 1 and the supplementary video.
-These experiments were conducted around a university campus, including pedestrian walkways, wooden bridges, grass-covered hills, beach volleyball courts, and gravel paths.
-Extensive outdoor testing highlights the policy’s capability to maintain a reliable and natural walking gait and withstand diverse environmental conditions, such as walking on wet surfaces and operating with strong wind.
+These experiments were conducted around a university campus, including walkways, wooden bridges, grass hills, beach volleyball courts, and gravel paths.
 Surprisingly, the policy shows generalization ability to scenarios not included in the training.
@@
 Despite the robot not being calibrated or trained for such conditions, the policy adapts to the environmental changes without additional training.
-This demonstrates the robustness and adaptability of our approach.
 Next, we examine the policy’s performance on terrains with obstacles.
 The first test involves a crate of gravel shown in Fig.
-6(b).
+7(b).
 The robot consistently performs stepping-in-place actions with a stable motion.
@@
 In the second test, we deploy the robot on a slippery terrain, where we distribute poppy seeds on a whiteboard shown in Fig.
-6(c).
+1.
 Surprisingly, our robot does not exhibit any perceptible shaking motions.
 In contrast, the company controller provided by Agility Robotics fails to walk over.
-Furthermore, the policy enables omnidirectional control of the robot.
-Experiments with velocity commands demonstrate the robot’s directional control, which is crucial for practical deployment.
-Our policy can sustain some deformation and irregularity but may drift and lose balance on such occasions.
-Readers can find these experiments in the video.
+Figure 7: (a) we perturb the robot with a harness. (b) we test the step-in-place motion on a crate of gravel. We let the robot walk over (c) high stairs (77cm + 99cm). (d) hill with gravel.
 VI-B Perturbation experiments
-We evaluate the policy’s response to external perturbations applied during locomotion.
+We evaluate the policy’s response to external perturbations during locomotion.
 Two scenarios are considered: pushing the robot’s center of mass (CoM) from the front, and back using a stick.
 The policy demonstrates robustness to withstand frontal and rearward pushes while maintaining a stable walking gait.
-We could not push the robot to failure even when applying a large force.
-Additionally, we apply a more substantial perturbation by using a harness to pull the robot with an impulsive force (see Fig.
-6(a)).
-The robot adapts dynamically, exhibiting agile adjustments to its gait to compensate for the external perturbations.
+Additionally, we apply a more substantial perturbation using a harness to pull the robot with an impulsive force (see Fig.
+7(a)).
+The robot adapts dynamically, exhibiting agile adjustments to compensate for the pulling.
 VII Conclusion
 We introduced L2T-RL, a novel single-stage learning framework that unifies teacher and student training to address sample inefficiency and enhance real-world performance.
-Our extensive simulation and hardware experiments demonstrate that L2T-RL achieves robust, agile, and precise locomotion while reducing sample complexity by $50\%$ and dramatically saving training time.
+Our extensive simulation and hardware experiments demonstrate that L2T-RL achieves robust, and agile locomotion while reducing sample complexity by $50\%$ and thus dramatically saving training time.
 These results highlight our contributions to redefining teacher-student learning paradigms and paving the way for practical RL-based robotic systems.